Alyeska Services Company, a division of a major oil company, provides various services to the operators of the North Slope oil field in Alaska. Data concerning the most recent year appear below:


Sales $7,500,000
Net operating income $600,000
Average operating assets $5,000,000


1.
value:
10 points


Requirement 1:
Compute the margin for Alyeska Services Company. (Omit the "%" sign in your response.)

Margin %

eBook Linkreferences


2.
value:
10 points

Requirement 2:
Compute the turnover for Alyeska Services Company. (Round your answer to 2 decimal places.)

Turnover

eBook Linkreferences


3.
value:
10 points


Requirement 3:
Compute the return on investment (ROI) for Alyeska Services Company. (Omit the "%" sign in your response.)

ROI %
Answer #1
Concepts and reason

Margin: It refers to the proportion of amount earned by the company over the sales made by the company during the respective financial year. Generally, the margin is expressed in terms of %.

Operating assets Turnover ratio: It refers to the ratio computed by dividing the sales made by the company during the year with the average operating assets.

Return on Investments: It refers to the amount of income earned by the company on the proportion of investments made by them during the year.

Fundamentals

Net operating income: It refers to the income earned by the company from its operations during the year. It can be computed by deducting the non-operating income from the total income of the company.

Sales: It refers to the revenue earned by the company by selling the goods or by rendering the service to the customers during a respective year.

Average operating assets: It refers to the average of the opening assets and the closing assets of the company. It can be said as the average amount invested by the company in the business on any particular date.

1.

Compute the margin for the company A by substituting the values as mentioned below-

Margin=NetoperatingIncomeSales=$600,000$7,500,000=0.08\begin{array}{c}\\{\rm{Margin}} = \frac{{{\rm{Net operating Income}}}}{{{\rm{Sales}}}}\\\\ = \frac{{\$ 600,000}}{{\$ 7,500,000}}\\\\ = 0.08\\\end{array}

The margin for the company S is 8%.

2.

Compute the operating assets turnover of the company A as mentioned below-

Turnover=SalesAverageoperatingassets=$7,500,000$5,000,000=1.50\begin{array}{c}\\{\rm{Turnover}} = \frac{{{\rm{Sales}}}}{{{\rm{Average operating assets}}}}\\\\ = \frac{{\$ 7,500,000}}{{\$ 5,000,000}}\\\\ = 1.50\\\end{array}

Turnover of the company is 1.50

3.

Compute the return on investments of the company A as mentioned below-

ReturnonInvestments=NetoperatingIncomeAverageoperatingAssets=$600,000$5,000,000=12.0\begin{array}{c}\\{\rm{Return on Investments}} = \frac{{{\rm{Net operating Income}}}}{{{\rm{Average operating Assets}}}}\\\\ = \frac{{\$ 600,000}}{{\$ 5,000,000}}\\\\ = 12.0\\\end{array}

% Return on investments is 12.

Ans: Part 1

Margin of the company is 8%

Part 2

Turnover of the company is 1.50

Part 3

% Return on investments is 12.